Integral of sin(x^2)


1
Expert's answer
2021-11-17T12:01:28-0500

"\smallint{sin(x^2)}.dx"



Substitute "u=\frac{\sqrt{2}.x}{\sqrt{π}}"


"dx=\frac{\sqrt{π}}{\sqrt{2}}.du"


"\smallint{sin(x^2)}.dx=
\frac{\sqrt{π}}{\sqrt{2}}\smallint{sin(\frac{π.u^2}{2})}.du"


These is special integral Fresnel integral "=S(u)"


Plug in solved integrals


"\frac{\sqrt{π}}{\sqrt{2}}\smallint{sin(\frac{π.u^2}{2})}.du=\frac{\sqrt{π}.S(u)}{\sqrt2}"



Undo Substitution "u=\frac{\sqrt{2}.x}{\sqrt{π}}"


"=\frac{\sqrt{π}.S(
\frac{\sqrt{2}.x}{\sqrt{π}}

)}{\sqrt2}+c"


Hence

"\smallint{sin(x^2)}.dx
=\frac{\sqrt{π}.S(
\frac{\sqrt{2}.x}{\sqrt{π}}

)}{\sqrt2}+c"